### v1

# Conformal online model aggregation

Abstract
Conformal prediction equips machine learning models with a reasonable notion of uncertainty quantification without making strong distributional assumptions. It wraps around any black-box prediction model and converts point predictions into set predictions that have a predefined marginal coverage guarantee. However, conformal prediction only works if we fix the underlying machine learning model in advance. A relatively unaddressed issue in conformal prediction is that of model selection and/or aggregation: for a given problem, which of the plethora of prediction methods (random forests, neural nets, regularized linear models, etc.) should we conformalize? This paper proposes a new approach towards conformal model aggregation in online settings that is based on combining the prediction sets from several algorithms by voting, where weights on the models are adapted over time based on past performance.
1 Introduction
Conformal prediction has emerged as a significant approach within the field of machine learning, offering a reliable procedure for uncertainty quantification. Unlike traditional methods that rely on strict distributional assumptions, conformal prediction provides a flexible framework by transforming point predictions obtained with any black-box model into confidence sets with valid finite-sample marginal coverage. Initially introduced by Vovk et al., (2005), conformal prediction has gained increasing prominence and has been extended beyond the i.i.d. (or exchangeable) data setting, some examples are Tibshirani et al., (2019); Barber et al., (2023); Gibbs and Candès, (2021).
A significant challenge in the application of conformal prediction arises in the context of model selection or aggregation. In fact, the conventional approach involves fixing the underlying machine learning (or prediction) model in advance and then obtaining an interval with a pre-specified level of marginal coverage. Furthermore, since with different models it is possible to obtain sets with the same theoretical guarantees, it becomes difficult to assess which model is the most suitable.
In this work, we propose a novel strategy based on the combination of prediction sets from multiple algorithms through a voting mechanism. Importantly, the weights assigned to each model in the aggregation process are dynamically adjusted over time, reflecting the past performance of the models. The performance of the various models will be assessed based on the size of the generated intervals. Specifically, models consistently producing smaller intervals will be favored over others. In fact, from a statistical point of view, Casella et al., (1993) outlines the fact that two quantities are important to evaluate a set: the level of coverage and the size. In the realm of conformal prediction, these two concepts are translated as validity and efficiency (Shafer and Vovk,, 2008). This implies that if the starting intervals are valid, then they will be evaluated for their efficiency.
1.1 Problem setup
We focus our attention on an online setting in which data arrive sequentially over time. A plausible scenario arises when dealing with a continuous flow of data; in symbols, for each iteration $t=1,2,...,T$, the covariate-response pair $z^{(t)}=(x^{(t)},y^{(t)})\in\mathcal{X}\times\mathcal{Y}\subseteq\mathbb{R}^{d}\times\mathbb{R}$ is observed.
In our scenario, we assume the presence of $K$ different prediction algorithms (or experts), denoted as $\hat{\mu}_{k}$, used to predict $y^{(t)}$ based on a given set of covariates $x^{(t)}$. As a result, in each iteration, we can derive $K$ distinct conformal prediction intervals (one for each model) $\mathcal{C}_{1}^{(t)},\dots,\mathcal{C}_{K}^{(t)}$ using the information from the set $\{z^{(i)}\}_{i=1}^{t-1}$. Actually, one can choose to use only a subset of $\{z^{(i)}\}_{i=1}^{t-1}$ to produce the interval due to computational issues or distribution shifts. If the data exhibit exchangeability, these sets have valid marginal coverage. This implies that, for any $\alpha\in(0,1)$, we achieve:
$\mathbb{P}\left(Y^{(t)}\in\mathcal{C}_{k}^{(t)}\right)\geq 1-\alpha,\quad k=1,\dots,K,$ (1)
where $\alpha$ is a user-defined error level and the probability is taken with respect to the new data point and the training points.
In light of the fact that our method is designed to return a single set from a collection of intervals, the initial question that arises is how to merge efficiently different conformal prediction intervals. It is crucial to highlight that these intervals are not independent, as they are constructed using the same data. One potential solution is the approach proposed by Gasparin and Ramdas, (2024), which involves a weighted majority vote procedure. Let $w^{(t)}=(w^{(t)}_{1},\dots,w^{(t)}_{K})$ be a set of weights, such that
$w^{(t)}_{k}\in[0,1],\quad k=1,\dots,K,$ (2)
$\sum_{k=1}^{K}w^{(t)}_{k}=1,$ (3)
where the weight $w_{k}^{(t)}$ represent the importance of the expert $k$ at the time $t$ in the voting procedure. We can define the majority vote set as
$\mathcal{C}^{(t)}:=\left\{y\in\mathcal{Y}:\sum_{k=1}^{K}w_{k}^{(t)}\mathds{1}\left\{y\in\mathcal{C}_{k}^{(t)}\right\}>\frac{1+u}{2}\right\},$ (4)
where $u$ is a realization from a uniformly distributed random variable in the interval $[0,1]$. If the weights are chosen a priori, it is possible to prove that if the starting intervals respect the property (1) then $\mathcal{C}^{(t)}$ has a valid $1-2\alpha$ coverage. In light of the selection of weights based on the data, they themselves become random quantities. Specifically, we have that $w^{(t)}$ is a realization of a random vector $W^{(t)}$ mapping the first $t-1$ observations into the $(K-1)$-dimensional simplex. Formally, we have $W^{(t)}:(\mathcal{X}\times\mathcal{Y})^{t-1}\to[0,1]^{K}$ with the additional constraint that the weights have to sum to one. In the following, we define the miscoverage indicator for the $k$-th interval at time $t$ as $\phi_{k}^{(t)}:=\mathds{1}\{y^{(t)}\notin\mathcal{C}_{k}^{(t)}\}$.
Proposition 1.1.
Let $\mathcal{C}_{1}^{(t)},\dots,\mathcal{C}_{K}^{(t)}$ be $K\geq 2$ different conformal prediction intervals that satisfy the property in (1). Let $W^{(t)}$ be a random vector having values in the $(K-1)$-dimensional simplex. Under the assumption that $\phi_{k}^{(t)}$ and $W_{k}^{(t)}$ are independent or negatively associated, for all $k=1,\dots,K$, then the set defined in (4) is such that $\mathbb{P}(Y^{(t)}\in\mathcal{C}^{(t)})\geq 1-2\alpha$.
Proof.
Due to the property described in (1) we have that $\mathbb{E}[\phi_{k}^{(t)}]\leq\alpha$, $k=1,\dots,K$. Now we can write
$\begin{split}\mathbb{P}\left(Y^{(t)}\notin\mathcal{C}^{(t)}\right)&=\mathbb{P}\left(\sum_{k=1}^{K}W_{k}^{(t)}\phi_{k}^{(t)}\geq\frac{1}{2}-U/2\right)\stackrel{{\scriptstyle(i)}}{{\leq}}2\mathbb{E}\left[\sum_{k=1}^{K}W_{k}^{(t)}\phi_{k}^{(t)}\right]\\ &\stackrel{{\scriptstyle(ii)}}{{\leq}}2\sum_{k=1}^{K}\mathbb{E}[W_{k}^{(t)}]\,\mathbb{E}[\phi_{k}^{(t)}]\leq 2\alpha\sum_{k=1}^{K}\mathbb{E}[W_{k}^{(t)}]=2\alpha,\end{split}$
where $(i)$ is valid due to the randomized Markov’s inequality described in Ramdas and Manole, (2023), while $(ii)$ holds due to the assumption regarding the independence or the negative association between the variables $\phi_{k}^{(t)}$ and $W_{k}^{(t)}$. In particular, $(ii)$ holds with equality under independence. ∎
The case of independence arises when the data are i.i.d. and the $K$ models are pre-trained and not iteratively updated or, more generally, when the models are trained in a different dataset than the one used for weights.
We remark the fact that if randomization is not allowed, then one can set the value of $u$ to 00 and the coverage guarantee still holds since $u>0$ almost surely. In addition, we highlight that if the weight vector takes the form $w^{(t)}_{k}=1$, and $w^{(t)}_{j}=0$ for all $j\neq k$, then our procedure reduces to the selection of the $k$-th model. A potential drawback of the procedure is that, even when initiated with a collection of intervals, it may produce a union of intervals. This does not occur if the intersection of the initial intervals is non-empty.
Consistent with our previous discussion, the loss of the different models is measured through the size of the outputted intervals. We define the loss function $\ell(\cdot)$ as
$\ell:=\ell(\mathcal{C})=\text{Lebesgue measure}(\mathcal{C}),$ (5)
where $\mathcal{C}$ represent an arbitrary set. In general, one could choose some nondecreasing function of the Lebesgue measure, but the identity seems to be a sensible canonical choice, so we stick to it in this paper. At each iteration, we observe losses of the experts $\ell^{(t)}=(\ell_{1}^{(t)},\dots,\ell_{K}^{(t)})\in\{\mathbb{R}^{+}\}^{K}$ and the cumulative loss for a given method after $t$ rounds is given by $L_{k}^{(t)}=\ell^{(1)}_{k}+\dots+\ell^{(t)}_{k}$.
A notable property of the majority vote is that it generates sets that are never larger than twice the weighted average size of the initial sets. For completeness, we present here an useful Lemma, which will be used in the following.
Lemma 1.2.
Let $\mathcal{C}^{(t)}$ be the set defined in (4), $\ell_{M}^{(t)}=\ell(\mathcal{C}^{(t)})$ be the Lebesgue measure associated to $\mathcal{C}^{(t)}$ and $w^{(t)}$ be a realization of $W^{(t)}$. Then,
$\ell_{M}^{(t)}=\ell(\mathcal{C}^{(t)})\leq 2\sum_{k=1}^{K}w_{k}^{(t)}\ell(\mathcal{C}_{k}^{(t)}).$ (6)
The proof is given in Gasparin and Ramdas, (2024).
From equation (6), it is now clear that our goal is to find a good weighting system in order to narrow down the size of $\mathcal{C}^{(t)}$, particularly we aim to assign a higher weight to experts that consistently yield smaller intervals compared to others.
In the following, we describe a general methodology to infer the weights assigned to the various experts over time. In particular, we start by describing the method under an i.i.d. setting and then extend the approach under a (possible) continuous distribution shift over time, as described in Gibbs and Candès, (2021). We will then demonstrate the practical effectiveness of the method through a series of simulations and examples using real-world data.
2 Related work
The so-called full conformal prediction was initially introduced by Vovk et al., (2005), while more computationally efficient versions were proposed by Papadopoulos et al., (2002), Lei et al., (2018) and Romano et al., (2019). In particular, the latter methods propose dividing the training set into two parts: a proper training set and a calibration set. All these approaches provide reliable uncertainty quantification for a single black-box algorithm and are grounded on the assumption that observations are independent and identically distributed or at least exchangeable.
In recent years, numerous works have aimed to extend conformal prediction beyond the context of i.i.d. or exchangeable data (Barber et al.,, 2023; Tibshirani et al.,, 2019). In particular, some of these studies have specialized in the case where data arrive in an online fashion and can exhibit a continuous shift in the distribution. One of the pioneering works is Gibbs and Candès, (2021), where the main idea is to adapt the error level $\alpha$ over time in order to achieve the desired error rate. Several works follow this path and extend this method in different ways (Gibbs and Candès,, 2023; Angelopoulos et al., 2024a, ; Zaffran et al.,, 2022; Angelopoulos et al., 2024b, ). Online learning plays a key role in this framework; indeed, the proposed methods try to align over time the miscoverage level (or a prespecified quantile) according to the most recent observations.
Lastly, outside the conformal framework, online learning is used to develop sequential prediction algorithms when multiple agents construct a prediction at each time. In this particular setting, it is customary that at each iteration a weight is assigned to the agent, which is subsequently updated in accordance with its performance. One of the most widely employed methods is the Exponential Weighted (EW) algorithm (Freund and Schapire,, 1997), which updates the weights in an inversely proportional manner to the exponential of the loss incurred at the $t$-th iteration. One of its extensions is proposed in de Rooij et al., (2014), and it has been demonstrated to be effective even in scenarios where the data is not adversarial.
3 Dynamic merging via exponential weighted majority
As illustrated in Section 1.1, our goal is to achieve a set that is as small as possible, particularly (6) asserts that this can be accomplished by learning a weight system that prioritizes methods yielding smaller intervals. Furthermore, we want that the cumulative loss of our proposed method, indicated by $L_{M}^{(t)}=\ell_{M}^{(1)}+\dots+\ell_{M}^{(t)}$, is not too large if compared to the cumulative loss incurred by the best expert up to time, indicated by
$L_{*}^{(t)}:=\min_{k\in\{1,\dots,K\}}L_{k}^{(t)}.$
A possible way to update our vector of weights is the Exponential Weight (ew) or Hedge Algorithm introduced in Freund and Schapire, (1997), which corresponds to Algortihm 1 with $\eta^{(t)}=\eta$, for all $t$. One way to evaluate the weighted majority method is the hedge loss $H^{(t)}=h^{(1)}+\dots+h^{(t)}$, where $h^{(t)}$ is the dot product $h^{(t)}=w^{(t)}\cdot\ell^{(t)}$. This can be interpreted as the expected loss achieved by randomly selecting the expert $k$ with probability $w_{k}^{(t)}$.
Data: $\mathcal{C}_{1}^{(t)},\dots,\mathcal{C}_{K}^{(t)}$ at each round $t$, initial learning rate $\eta^{(0)}\geq 0$
$w_{k}^{(1)}\leftarrow 1/K,L_{k}^{(0)}\leftarrow 0,$ $k=1,\dots,K$;
for rounds $t=1,\dots,T$ do do
       $\mathcal{C}^{(t)}\leftarrow\left\{s\in\mathcal{S}:\sum_{k=1}^{K}w_{k}^{(t)}\mathds{1}\{s\in\mathcal{C}_{k}^{(t)}\}>\frac{1}{2}\right\}$;
       Receive loss $\ell_{k}^{(t)}$, update $L_{k}^{(t)}:=L_{k}^{(t-1)}+\ell_{k}^{(t)}$;
       Update learning rate $\eta^{(t)}$;
       $w_{k}^{(t+1)}\leftarrow\exp\{-\eta^{(t)}L_{k}^{(t)}\}/\sum_{j=1}^{K}\exp\{-\eta^{(t)}L_{j}^{(t)}\}$;
end for
Algorithm 1 Exponentially Weighted Majority Vote
The algorithm updates the weights according to the performance of the methods during the different rounds inversely proportional to the exponential of their size multiplied by a parameter $\eta$. This “learning rate” $\eta$ plays a crucial role; if $\eta$ approaches zero, the weights approach a uniform weighting, and if $\eta\to\infty$ the algorithm reduces to the Follow-the-Leader strategy, which puts all the weight on the method with the smallest loss so far, as explained in de Rooij et al., (2014). In certain scenarios, determining the appropriate value for this parameter may be difficult; indeed, we do not know if there is a clear method outperforming the others or any constant works well in some situations but not in others.
Our proposed solution is to employ the AdaHedge algorithm, which dynamically adjusts the learning parameter over time as follows. The weights at each round are assigned according to $w^{(t)}_{k}\propto\exp\{-\eta^{(t)}L^{(t-1)}_{k}\}$, $k=1,\dots,K$, with
$\eta^{(t)}:=\frac{\ln K}{\delta^{(1)}+\dots+\delta^{(t-1)}},$
where $\delta^{(t)}=h^{(t)}-m^{(t)}$ is the difference between the hedge loss $h^{(t)}$ and the mix loss,
$m^{(t)}:=-\frac{1}{\eta^{(t)}}\ln\left(w^{(t)}\cdot\exp\{-\eta^{(t)}\ell^{(t)}\}\right).$
The AdaHedge algorithm provides an upper bound on regret, defined as the difference between the Hedge loss and the loss of the best method up to time $t$, specifically:
$\mathcal{R}^{(t)}=H^{(t)}-L_{*}^{(t)}.$
Consider $\ell_{+}^{(t)}=\max_{k}\ell_{k}^{(t)}$ and $\ell_{-}^{(t)}=\min_{k}\ell_{k}^{(t)}$ to represent the smallest and largest loss in round $t$, and let $L_{+}^{(t)}$ and $L_{-}^{(t)}$ represent their cumulative sum. Additionally, define $S^{(t)}=\max\{s^{(1)},\dots,s^{(t)}\}$ denote the largest loss range, where $s^{(t)}=l_{+}^{(t)}-l_{-}^{(t)}$, after $t$ trials. Suppose our objective is to find an upper bound for regret after an arbitrary number of rounds $T$. According to Theorem 8 in de Rooij et al., (2014), we have
$H^{(T)}\leq L_{*}^{(T)}+2\left(S^{(T)}\ln K\frac{(L_{+}^{(T)}-L_{*}^{(T)})(L_{*}^{(T)}-L_{-}^{(T)})}{L_{+}^{(T)}-L_{-}^{(T)}}\right)^{1/2}+S^{(T)}\left(\frac{16}{3}\ln K+2\right).$ (7)
It should be noted that this upper limit depends on the observed losses but not on the number of trials. This implies that the bound remains unchanged when additional rounds where all experts have the same loss are introduced. Another noteworthy feature of the method is its invariance to translations or scalings of the loss function. Specifically, the sequence of weights remains unchanged when the loss vectors $\ell^{(t)}$ are multiplied by a positive constant and (or) a quantity is added to them.
It is possible to bound the total measure of the sets produced by the majority vote method.
Proposition 3.1.
The total Lebesgue measure of the weighted majority vote set over $T$ rounds can be bounded as
$L_{M}^{(T)}\leq 2L_{*}^{(T)}+4\left(S^{(T)}\ln K\frac{(L_{+}^{(T)}-L_{*}^{(T)})(L_{*}^{(T)}-L_{-}^{(T)})}{L_{+}^{(T)}-L_{-}^{(T)}}\right)^{1/2}+2S^{(T)}\left(\frac{16}{3}\ln K+2\right).$ (8)
Proof.
From Lemma 1.2 we have that $\ell_{M}^{(t)}\leq 2h^{(t)}$. This implies that $L_{M}^{(t)}=\ell_{M}^{(1)}+\dots+\ell_{M}^{(t)}\leq 2H^{(t)}$; combining this result with (7) gives the desired bound. ∎
The above expression is identical to the regret bound (7), except for an additional factor of 2.
In certain instances, the Lebesgue measure of the set cannot be directly employed because of situations where the width of some intervals can be unbounded. For example, in adaptive conformal prediction (Gibbs and Candès,, 2021), there is no guarantee that the value of $\alpha$ is greater than zero. A possible solution is to define the loss function as an increasing, nonnegative, bounded function of the measure of the set (so that the weight of a predictor does not get permanently set to zero if its $\alpha$ happens to be negative in some round). We will discuss this example in the following sections.
4 Conformal online model aggregation: i.i.d. setting
Code to reproduce all experiments can be found at https://github.com/matteogaspa/coma.
As explained at the beginning, a relatively open question in the conformal prediction literature is how to do model selection or aggregation using conformal prediction: if we have $K$ different prediction algorithms that one could be using within conformal prediction, how do we combine their predictions so that we can essentially do as well as the “best” predictor? Since all methods do provide the same coverage level of $1-\alpha$ under i.i.d. data, we use a working definition of the “best” predictor as the one whose expected interval length is shortest.
Below, we provide one answer to this question in an online or streaming data setting, by employing the algorithms introduced in the previous section. When our dynamic merging Algorithm 1 is used (with constant, adaptive or another stepsize rule) in this conformal context, we title the method as “conformal online model aggregation”, or COMA for short. In the i.i.d. setting, it is often the case that our adaptive stepsize update rule results in the weight vector rapidly putting almost full (unit) weight on the best predictor, resulting effectively in online model “selection”, but experiments in non-i.i.d. settings suggest that the weight vector can meaningfully fluctuate between different predictors if they perform better in different periods.
We start by examining our proposed methods, in a scenario where data are i.i.d., this implies that each algorithm at each iteration provides an interval with the guarantee described in (1), while our procedure has a valid $1-2\alpha$ coverage level.
To examine the practical behavior of COMA, we apply it to a real-world dataset. Specifically, the dataset comprises $75000$ observations pertaining to Airbnb apartments in New York City. The response variable is represented by the logarithm of the nightly price, while the covariates include information about the apartment and its geographical location. The data has been partitioned into 75 rounds, each consisting of $1000$ observations. At each iteration, $K$ split-conformal prediction intervals are constructed using various algorithms. In the first simulation, $K=4$ regression algorithms were employed: linear model, lasso, ridge, and random forest. In the second scenario $K=5$ and also a neural network was utilized.
Figure 1: Figure shows the performance of our conformal online model aggregation algorithm. We plot the hedge loss ($h_{t}$) obtained during various iterations with either a constant or adaptive learning rate scheme, along with the loss of the Random Forest algorithm (left), and the loss from both Random Forest and Neural Network (right). The series have been smoothed using a moving average $(5,\frac{1}{5})$. In both cases, COMA with the adaptive learning rate schedule (and with rate equal to 1) quickly learns that RF is the best method.
The weights have been updated with the COMA with fixed or adaptive $\eta$ methods described above. For Algorithm 1 with variable learning parameter throughout the iterations, two different values were selected, $\eta=0.1$ and $\eta=1$. In Figure 1, the hedge loss during various iterations in both scenarios is presented, along with the loss obtained by the best performing method(s). It can be observed that in both cases, the hedge loss of the COMA algorithm with $\eta=1$ and of the adaptive COMA stabilizes at the level of loss obtained by the random forest after a certain number of iterations, which is the best method. On the other hand, if $\eta$ is set to 0.1 the loss exhibits a slower decrease, attributed to the lower value of the learning parameter. In Figure 6 and Figure 5 of Appendix A, we report the weights assumed by the different regression algorithms during certain iterations. It can be seen that the COMA with adaptive parameter, after a few iterations, concentrates all the mass in the random forest, while the COMA algorithm with fixed learning parameter is more conservative in assigning the weights. This fact can be explained by the plot in Figure 7 of Appendix A where it is possible to note that in both cases ($K=4$ and $K=5$) the learning parameter $\eta^{(t)}$ is always higher than the levels $\eta=1$ and $\eta=0.1$.
5 COMA under distribution shift
As explained in Section 1, conformal prediction offers a general approach to convert the output of any prediction algorithm into a prediction set. Although its validity hinges on the assumption of exchangeability, this assumption is frequently violated in numerous real-world applications. In the energy market, for instance, consumer behavior can undergo significant shifts in response to new legislation, major global events, and financial market fluctuations. In general, where the distribution of the data shift over time we cannot expect to obtain reliable results using standard conformal inference. In recent years, several works have been proposed to address this issue; some examples are Gibbs and Candès, (2021) and Angelopoulos et al., 2024a .
In Gibbs and Candès, (2021), adaptive conformal inference (ACI) is introduced as a novel method for constructing prediction sets in an online manner, providing robustness to changes in the marginal distribution of the data. Their work is based on the update of the error level $\alpha$ to achieve the desired level of confidence. Indeed, given the non-stationary nature of the data-generating distribution, conventional results do not guarantee a $1-\alpha$ coverage. However, at each time $t$, an alternative value $\alpha^{(t)}$ might exist, allowing the attainment of the desired coverage.
The procedure described in Gibbs and Candès, (2021) recursively updates the error rate $\alpha$ as follows:
$\alpha_{k}^{(1)}=\alpha,$ (9)
$\alpha_{k}^{(t)}=\alpha_{k}^{(t-1)}+\gamma(\alpha-\phi_{k}^{(t-1)}),\quad t\geq 2,$ (10)
where $\phi^{(t)}_{k}=\mathds{1}\{Y_{t}\notin\mathcal{C}_{k}^{(t)}(\alpha_{k}^{(t)})\}$ is the sequence of miscoverage events obtained by the $k$-th agent and setting the error to $\alpha_{k}^{(t)}$, while $\gamma>0$ represents a step size parameter. Possible improvements are proposed in Gibbs and Candès, (2023) and Zaffran et al., (2022), in particular, they tune the parameter $\gamma$ over time to make the procedure more flexible. A drawback arises from the fact that in certain situations, ACI can return either $\mathbb{R}$ or $\emptyset$ as a set, attributed to the circumstance that in some iterations, the parameter $\alpha^{(t)}$ can be smaller than 0 or greater than 1.
Another method is presented in Angelopoulos et al., 2024a , where instead of updating the error level $\alpha$, there is an update of the quantile of the distribution of scores. In this setting, we define $s^{(t)}:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ as the conformal score function, which measures the accuracy of the prediction at time $t$. Actually, for different models one can use different scores (e.g. absolute or signed residuals), but for clarity of exposition, we will use the same $s(\cdot,\cdot)$ for all experts. The procedure updates the quantiles as follows:
$q_{k}^{(1)}=q,$
$q_{k}^{(t)}=q_{k}^{(t-1)}+\gamma(\phi_{k}^{(t-1)}-\alpha),\quad t\geq 2,$
where $\phi_{k}^{(t-1)}=\mathds{1}\{y^{(t)}\notin\mathcal{C}_{k}^{(t)}(q_{k}^{(t)})\}$ and the sets at each time are simply defined by
$\mathcal{C}_{k}^{(t)}=\left\{y\in\mathcal{Y}:\,s^{(t)}(x^{(t)},y)\leq q_{k}^{(t)}\right\}.$
Also in this case some improvements are proposed, such as an adaptive learning rate or the introduction of an error integrator. The proposal in Angelopoulos et al., 2024b is to set a decaying step size $\gamma^{(t)}$. In particular, they prove that if scores are bounded and $\gamma^{(t)}\propto t^{-1/2-\epsilon}$ for some $\epsilon\in(0,1/2)$, then the miscoverage coverage error is bounded as $\mathcal{O}(\frac{1}{(T^{1/2}-\epsilon)})$ in the adversarial setting, while in the i.i.d. the algorithm achieves a coverage guarantee.
In light of the dynamics shift observed in the marginal distribution of the data, it is conceivable that the weights assigned to different algorithms may also undergo variations throughout iterations. We now describe two different ways to combine our dynamic ensembling algorithm with ACI: either as a wrapper that does not alter the inner workings of ACI (Subsection 5.1), or by altering the feedback provided to ACI itself (Subsection 5.2).
5.1 Decentralized COMA under distribution shift
In the scenario described above, suppose that each agent operates independently by providing an interval at each iteration and, subsequently, the aggregator merges these intervals according to the vector $w^{(t)}$. This framework can be referred to as a decentralized dynamic merging. Under this setting, the advantage of the quantile tracking method is that it never returns sets of infinite dimensions. This implies that the loss function used can remain the size of the different sets, ensuring that the bound described in (8) is valid. In the case of using the described ACI approach, it becomes necessary to employ a function $g(\cdot)$ that is capable of returning a finite number when the miscoverage rate $\alpha_{k}^{(t)}$ is less than or equal to 0.
We use the methods described in Section 3 on adaptive conformal inference in a real data set. We use $K=3$ different regression algorithms, and we employ quantile tracking with a decaying step size $\gamma^{(t)}\propto t^{-1/2-\epsilon}$. In particular, we use the ELEC2 (Harries,, 2003) data set that monitors electricity consumption and pricing in the states of New South Wales and Victoria in Australia, with data recorded every 30 minutes over a 2.5-year period from 1996 to 1999. For our experiment, we utilize four covariates: nswprice and vicprice, representing the electricity prices in each respective state, and nswdemand and vicdemand, denoting the usage demand in each state. The response variable is transfer, indicating the quantity of electricity transferred between the two states. We narrow our focus to a subset of the data, retaining only observations within the time range of 9:00 am to 12:00 pm to mitigate daily fluctuation effects, the same procedure is chosen in Barber et al., (2023). Additionally, we discard an initial time segment during which the transfer value remains constant. The models used correspond to three distinct linear models based on different covariates. The first model relies on the entire set of covariates at time $t$, the second model employs $y^{(t-1)}$ and $y^{(t-2)}$ as covariates (it corresponds to an AR(2) without constraints on the parameters), while the third model supplements the latter with the covariate nwsprice.
Figure 2: The top row displays local coverage for adaptive conformal inference using quantile tracking for the three separate regression algorithms. The bottom row shows local coverage for our merged set using COMA with adaptive $\eta$ or $\eta=0.1$ (with and without randomization). In the bottom-left plot the lines of the randomized and non-randomized methods overlap as the weights concentrate after a few iterations onto a single model.
The performance of the methods is compared using the local level coverage on $200$ data points:
$\textrm{localCov}^{(t)}:=\sum_{i=t-100+1}^{t+100}\mathds{1}\left\{y^{(i)}\in\mathcal{C}^{(i)}\left(q_{k}^{(i)}\right)\right\},$ (11)
where $\mathcal{C}_{k}^{(t)}(q_{k}^{(t)})$ is the set obtained using the $k$-th regression algorithms with the quantile fixed at $q_{k}^{(t)}$. The local coverage is also used to measure the performance of our merging procedure. In particular, we compare the use of an adaptive learning parameter and a small fixed $\eta=0.1$. As can be seen in Figure 2, the COMA procedure with adaptive $\eta^{(t)}$ obtains a coverage that is similar to the prespecified level $1-\alpha$. This is because the weights concentrate rapidly on one model over time; in particular, this model is not always the same, as is possible from Figure 3. However, in both cases, the coverage is similar to the pre-specified level $1-\alpha$.
Figure 3: Weights (first row) and losses, $\ell^{(t)}$, (second row) obtained by the various model for the decentralized COMA algorithm applied in ELEC2 dataset. The series of the losses have been smoothed using a moving average $(100,\frac{1}{100})$. The loss of the weighted majority vote stabilizes around the two best models after few iterations. The strategy with an adaptive $\eta^{(t)}$ is more aggressive than the one with $\eta^{(t)}=\eta$.
5.2 Adaptive conformal inference directly applied on dynamic merging
The approach described in the preceding section had agents updating the quantiles of their score function based on their own past errors. This makes sense when all agents can observe the ground truth and calculate their own errors, and they do not care about the aggregator’s goals. In other words, the dynamic merging algorithm was just an outer wrapper that did not interfere with the inner functioning of the $K$ adaptive conformal inference algorithms.
Below, we show that our dynamic ensembling method can provide direct feedback to the adaptive conformal inference framework, and this can have some advantages in settings where the end goal is good aggregator performance only, and individual agents do no have their own goals (of maintaining coverage, say). First, we observe that if the absolute residuals are employed as the score function and if the quantile ($q^{(t)})$ is common across all algorithms, then the size of the sets at time $t$ is equal for all intervals. This implies that the weights are not updated during the various iterations. In this case, we propose to use a value $\alpha^{(t)}$ that is common for all $K$ agents, and it is updated by the aggregator based on the performance of the exponential weighted majority vote. In other words, the aggregator, at each iteration, tries to learn the coverage level required to obtain a $(1-\alpha)$-confidence set. The procedure recursively updates the $\alpha$ level according to the miscoverage event $\phi^{(t)}=\mathds{1}\{Y_{t}\notin\mathcal{C}^{(t)}(\alpha^{(t)})\}$, where
$\mathcal{C}^{(t)}:=\left\{y\in\mathbb{R}:\sum_{k=1}^{K}w_{k}^{(t)}\mathds{1}\left\{y\in\mathcal{C}_{k}^{(t)}\left(\alpha^{(t)}\right)\right\}>1/2\right\},$
and $w_{k}^{(t)}$ are the weights learned by the COMA procedure. The error level (common among the agents) is updated as in (9) and (10). It is important to remark that the length of the intervals can be different across the various agents also if $\alpha^{(t)}$ is common. In such a scenario, ACI operates directly on the set produced by the aggregator and so we directly recover the properties of the ACI method. Further, there is no need for randomization to enhance the length and coverage of the intervals. In particular, Gibbs and Candès, (2021) proves that the empirical error converges almost surely to $\alpha$ under weak conditions.
Figure 4: Local level coverage, weights assumed during the iterations and losses for direct ACI for COMA with adaptive $\eta$ (top) and $\eta=0.1$ (bottom) applied in ELEC2 dataset. The series of the losses have been smoothed using a moving average $(100,\frac{1}{100})$. In both cases, the local level coverage oscillates around the level $1-\alpha$.
We applied the procedure to the same dataset used previously, with the target coverage level fixed at 0.05 and the loss function is $\arctan(\ell^{(t)})$. The use of a function which bounds our loss is necessary since ACI can output $\mathbb{R}$ as interval. Also in this we use (11) to compare the various methods. The local level coverage oscillates around the prespecified while the it is possible to see that the strategy with an adaptive $\eta$ is more aggressive and adapts rapidly the weights according to the observed loss.
6 Summary
Our work presents a method for aggregating conformal prediction intervals derived from different algorithms. Specifically, a method based on majority voting is proposed, where the weight of each individual agent is updated based on its past performance. In some cases, when there is an algorithm that outperforms the others, the weight is assigned to the better-performing model, resulting in model selection. Particularly, in the case of i.i.d. data, it is very common for there to exist a better model that outperforms the others in terms of size of the outputted sets.
We also tested our method in cases where the data underwent a distribution shift, along with quantile tracking (Angelopoulos et al., 2024a, ; Angelopoulos et al., 2024b, ) and adaptive conformal inference (Gibbs and Candès,, 2021) methods. Specifically, two contexts are explored: in the first case, each model operates independently, while in the second case, the aggregator acts above the models, and adaptive conformal inference operates directly on it.
Acknowledgments
AR acknowledges funding from NSF grant IIS-2229881, and thanks Rina Foygel Barber, Ryan Tibshirani and Emmanuel Candès for useful conversations around model selection in conformal prediction.
Appendix A Additional results from Section 4
Figure 5: Weights assumed by the regression algorithms during different iterations. The regression algorithms used are lasso, linear model, ridge and random forest.
Figure 6: Weights assumed by the regression algorithms during different iterations. The regression algorithms used are lasso, linear model, neural net, ridge and random forest.
Figure 7: Series of values taken by the learning parameter $\eta^{(t)}$ over time. The plot on the right corresponds to the plot on the left, zoomed in on the interval 0 to 10.
Appendix B Additional simulations on Adaptive Conformal Inference
Investigating the ACI (Adaptive Conformal Inference) method in conjunction with the adaptive COMA, we conduct a case study in which changes in marginal distribution also impact the weights of the two algorithms. We suppose to have two covariates, $x_{1}$ and $x_{2}$, and the data are generated in blocks, with only one of the two covariates influencing the response at a time. Specifically, at each time $t$ we have $y^{(t)}=2x^{(t)}+\epsilon^{(t)}$ where
$x^{(t)}=\begin{cases}x_{1}^{(t)},\quad t\in\{[0,50)\cup[100,150)\cup[250,350)\cup[450,550)\cup\dots\cup[1250,1350)\}\\ x_{2}^{(t)},\quad t\in\{[50,100)\cup[150,250)\cup[350,450)\cup\dots\cup[1350,1450]\}\end{cases}$
and $x_{1}^{(t)},x_{2}^{(t)},\epsilon^{(t)}$ are generated from a standard normal distribution. Two standard linear models are used, one using $x_{1}$ as a regressor and the other using $x_{2}$. Prediction intervals are generated using split conformal on the most recent 100 observations, and the parameter $\gamma$ for the Adaptive Conformal algorithm is set to 0.005. In addition, the cumulative density function of a $\textrm{Gamma}(1,0.1)$ is used to scale the length of the intervals. The entire procedure is repeated 2000 times. The choice of the weight assigned to each method dynamically adapts to the performance of the two approaches, evaluated in terms of cumulative loss, as can be seen in Figure 8. In particular, the weight assigned to one of the two experts is higher when the cumulative loss is lower than that of the other expert.
Figure 8: The mean of cumulative losses ($L_{1}^{(t)},L_{2}^{(t)}$) across 2000 replications throughout the iterations is depicted in the left plot for the linear model employing $x_{1}$ (orange) or $x_{2}$ (green) as the regressor. In the right plot, the average weights for 2000 replications are illustrated during the iterations of the linear model using either $x_{1}$ or $x_{2}$ as the regressor. In the background of both plots, the shaded region indicates iterations when $L_{1}^{(t)}>L_{2}^{(t)}$.