### v2

# Conformal online model aggregation

Abstract
Conformal prediction equips machine learning models with a reasonable notion of uncertainty quantification without making strong distributional assumptions. It wraps around any black-box prediction model and converts point predictions into set predictions that have a predefined marginal coverage guarantee. However, conformal prediction only works if we fix the underlying machine learning model in advance. A relatively unaddressed issue in conformal prediction is that of model selection and/or aggregation: for a given problem, which of the plethora of prediction methods (random forests, neural nets, regularized linear models, etc.) should we conformalize? This paper proposes a new approach towards conformal model aggregation in online settings that is based on combining the prediction sets from several algorithms by voting, where weights on the models are adapted over time based on past performance.
1 Introduction
Conformal prediction has emerged as a significant approach within the field of machine learning, offering a reliable procedure for uncertainty quantification. Unlike traditional methods that rely on strict distributional assumptions, conformal prediction provides a flexible framework by transforming point predictions obtained with any black-box model into prediction sets with valid finite-sample marginal coverage. Initially introduced by Vovk et al., (2005), conformal prediction has gained increasing prominence and has been extended beyond the i.i.d. (or exchangeable) data setting, some examples are Tibshirani et al., (2019); Barber et al., (2023); Gibbs and Candès, (2021).
A significant challenge in the application of conformal prediction arises in the context of model selection or aggregation. The conventional approach involves fixing the underlying machine learning (or prediction) model in advance and then obtaining an interval with a pre-specified level of marginal coverage. Since different models yield sets with the same theoretical coverage guarantees, the natural goal is to select the best, or perform as well as the best, of a collection of models. In this work, “best” refers to the set with smallest size, such as cardinality for a discrete set, or Lebesgue measure for a continuous set.
In this work, we propose a novel strategy based on the combination of prediction sets from multiple algorithms through a voting mechanism. Importantly, the weights assigned to each model in the aggregation process are dynamically adjusted over time, reflecting the past performance of the models. The performance of the various models will be assessed based on the size of the generated intervals. Specifically, models consistently producing smaller intervals will be favored over others. In fact, from a statistical point of view, Casella et al., (1993) outlines the fact that two quantities are important to evaluate a set: the level of coverage and the size. In the realm of conformal prediction, these two concepts are translated as validity and efficiency (Shafer and Vovk,, 2008). This implies that if the starting intervals are valid, then they will be evaluated for their efficiency.
1.1 Problem setup
We focus our attention on an online setting in which data arrive sequentially over time. A plausible scenario arises when dealing with a continuous flow of data; in symbols, for each iteration $t=1,2,...,T$, a covariate-response pair $z^{(t)}=(x^{(t)},y^{(t)})\in\mathcal{X}\times\mathcal{Y}$ representing an instance from the random vector $Z^{(t)}=(X^{(t)},Y^{(t)})$ is observed.
In our scenario, we assume the presence of $K$ different prediction algorithms (or experts), denoted as $\hat{\mu}_{k}$, used to predict $Y^{(t)}$ based on a given set of covariates $X^{(t)}$. As a result, in each iteration, we can derive $K$ distinct conformal prediction sets (one for each model) $\mathcal{C}_{1}^{(t)},\dots,\mathcal{C}_{K}^{(t)}$ based on the information from ($(\{z^{(i)}\}_{i=1}^{t-1},x^{(t)})$. Actually, one can choose to use only a subset of $\{z^{(i)}\}_{i=1}^{t-1}$ to produce the interval due to computational issues or distribution shifts. If the data exhibit exchangeability, these sets have valid marginal coverage. This implies that, for any $\alpha\in(0,1)$, we achieve:
$\mathbb{P}\left(Y^{(t)}\in\mathcal{C}_{k}^{(t)}\right)\geq 1-\alpha,\quad k=1,\dots,K,$ (1)
where $\alpha$ is a user-defined error level and the probability is taken with respect to the new data point and the training points.
In light of the fact that our method is designed to return a single set from a collection of intervals, the initial question that arises is how to efficiently merge different conformal prediction sets (i.e. conformal prediction sets produced by different underlying prediction algorithms). These intervals are not independent, as they are constructed using the same data. One potential solution is an approach recently proposed by Gasparin and Ramdas, (2024), which involves a weighted majority vote procedure. Let $w^{(t)}=(w^{(t)}_{1},\dots,w^{(t)}_{K})$ be a set of weights, such that
$w^{(t)}_{k}\in[0,1],\quad k=1,\dots,K,\quad\text{ and }\sum_{k=1}^{K}w^{(t)}_{k}=1,$ (2)
where the weight $w_{k}^{(t)}$ represent the importance of the expert $k$ at the time $t$ in the voting procedure. We can define the majority vote set as
$\mathcal{C}^{(t)}:=\left\{y\in\mathcal{Y}:\sum_{k=1}^{K}w_{k}^{(t)}\mathds{1}\left\{y\in\mathcal{C}_{k}^{(t)}\right\}>\frac{1+u^{(t)}}{2}\right\},$ (3)
where $u^{(t)}$ is a realization from a uniformly distributed random variable in the interval $[0,1]$ independent of all the data. If the weights are chosen a priori (independent of the data), Gasparin and Ramdas, (2024) prove that if the starting intervals are marginally valid (they respect property (1)) then $\mathcal{C}^{(t)}$ has $1-2\alpha$ coverage, and the factor 2 is unavoidable in general.
There are two hurdles to applying their ideas directly. First, in light of the selection of weights based on the data, the weights themselves become random quantities, so their theory does not immediately apply (especially for non-i.i.d. data). The second question is how one should update the weights in a theoretically sound way that performs well practically. The rest of this paper will present ways to overcome these hurdles.
1.2 Related work
The so-called full conformal prediction was initially introduced by Vovk et al., (2005), while more computationally efficient versions were proposed by Papadopoulos et al., (2002), Lei et al., (2018), Romano et al., (2019) and Romano et al., (2020). In particular, the latter methods propose dividing the training set into two parts: a proper training set and a calibration set. All these approaches provide reliable uncertainty quantification for a single black-box algorithm and are grounded on the assumption that observations are independent and identically distributed or at least exchangeable. Humbert et al., (2023) proposes a method for aggregating the results of various agents within a conformal prediction framework. This method specifically addresses the context of federated learning, where privacy concerns restrict agent operations. The problem of model selection, within the realm of conformal prediction, is studied by Yang and Kuchibhotla, (2021) and Fan et al., (2023). Our work is distinct from theirs, as it operates within an online setting and can be extended to scenarios where data are not independently and identically distributed. Furthermore, rather than performing model selection, our approach focuses on model aggregation.
In recent years, numerous works have aimed to extend conformal prediction beyond the context of i.i.d. or exchangeable data (Barber et al.,, 2023; Tibshirani et al.,, 2019). In particular, some of these studies have specialized in the case where data arrive in an online fashion and can exhibit a continuous shift in the distribution. One of the pioneering works is Gibbs and Candès, (2021), where the main idea is to adapt the error level $\alpha$ over time in order to achieve the desired error rate. Several works follow this path and extend this method in different ways (Gibbs and Candès,, 2023; Angelopoulos et al., 2024a, ; Zaffran et al.,, 2022; Angelopoulos et al., 2024b, ). Online learning plays a key role in this framework; indeed, the proposed methods try to align over time the miscoverage level (or a prespecified quantile) according to the most recent observations.
Lastly, outside the conformal framework, online learning is used to develop sequential prediction algorithms when multiple agents construct a prediction at each time. In this particular setting, it is customary that at each iteration a weight is assigned to the agent, which is subsequently updated in accordance with its performance. One of the most widely employed methods is the Exponential Weighted (EW) algorithm (Freund and Schapire,, 1997), which updates the weights in an inversely proportional manner to the exponential of the loss incurred at the $t$-th iteration. See Cesa-Bianchi and Lugosi, (2006) for an introduction of the method. One of its extensions is proposed in de Rooij et al., (2014), and it has been demonstrated to be effective even in scenarios where the data is not adversarial.
1.3 Summary of contributions
We introduce a method for updating the weights in (3) based on the performance obtained by the different models; where the performance is measured through the dimension of the set (cardinality if $\mathcal{Y}$ is a discrete set or size if $\mathcal{Y}\subseteq\mathbb{R}$). This method relies on the exponential weighted algorithm and enables the establishment of an upper bound on the observed loss in relation to the best-performing algorithm, after $T$ iterations. In addition, we provide the theoretical assumptions under which our method guarantees a valid prediction set of level $1-2\alpha$. We will then demonstrate the practical effectiveness of the method through a series of simulations and examples using real-world data.
2 Weighted majority with data-dependent weights
Suppose we have that $w^{(t)}$ is a realization of a random vector $W^{(t)}$ that maps the first $t-1$ observations of $\{Z^{(i)}\}_{i=1}^{t-1}$ into the $(K-1)$-dimensional simplex $\Delta_{K-1}:=\{w\in[0,1]^{K}:\mathbf{1}^{T}w=1\}$, or more succinctly, $W^{(t)}:(\mathcal{X}\times\mathcal{Y})^{t-1}\to\Delta_{K-1}$. In the following, we define the miscoverage indicator for the $k$-th set at time $t$ as $\phi_{k}^{(t)}:=\mathds{1}\{y^{(t)}\notin\mathcal{C}_{k}^{(t)}\}$ and $[K]:=\{1,\dots,K\}$. We will make one of the following assumptions:
  1. 1.
    Conditional coverage:
    $\mathbb{E}\left[\phi_{k}^{(t)}\mid\{Z^{(i)}\}_{i=1}^{t-1}\right]\leq\alpha,\text{ for all }k\in[K].$ (4)
    This holds, for example, for conformal prediction when the data are i.i.d.
  2. 2.
    Marginal coverage and negative total elementwise correlation between $\phi^{(t)}$, $W^{(t)}$:
    $\mathbb{E}\left[\phi_{k}^{(t)}\right]\leq\alpha\text{ for }k\in[K],\leavevmode\nobreak\ \text{ and }\mathbb{E}\left[\sum_{k=1}^{K}\left(\phi_{k}^{(t)}-\mathbb{E}[\phi_{k}^{(t)}]\right)\left(W_{k}^{(t)}-\mathbb{E}[W_{k}^{(t)}]\right)\right]\leq 0.$ (5)
    $\phi^{(t)}$, $W^{(t)}$ could be negatively correlated if, for example, it happens to be the case that predictors that were assigned high weights (based on past data) are less likely to make mistakes.
Theorem 2.1.
Let $\mathcal{C}_{1}^{(t)},\dots,\mathcal{C}_{K}^{(t)}$ be $K\geq 2$ different prediction intervals for $Y^{(t)}$, let $W^{(t)}$ be a random vector in $\Delta_{K-1}$ depending only on $\{Z^{(i)}\}_{i=1}^{t-1}$, and consider the weighted majority vote set $\mathcal{C}^{(t)}$ in (3). It always holds that $\mathbb{P}(Y^{(t)}\notin\mathcal{C}^{(t)})\leq 2\mathbb{E}[\max_{k}(\phi_{1}^{(t)},\dots,\phi^{(t)}_{K})]$. If either assumption (4) or (5) holds, then $\mathbb{P}(Y^{(t)}\notin\mathcal{C}^{(t)})\leq 2\alpha.$
The proof of the theorem is given in Appendix C.
Remark 2.2.
The first part of the theorem states, without any assumption, that the probability of miscoverage is bounded by $2\mathbb{E}[\max_{k}(\phi_{1}^{(t)},\dots,\phi^{(t)}_{K})]$. This can translate into a trivial bound (like 1) if there is at least one algorithm that always makes an error, while it can be approximately $2\alpha$ in the case where miscoverage events are strongly correlated and $\mathbb{E}[\phi^{(t)}_{k}]\leq\alpha$, for each $k=1,\dots,K$.
Remark 2.3.
An elementary instance where the negative total elementwise correlation stated in Assumption 2 in Theorem 2.1 is valid can be described as follows. Consider the scenario involving just two models, resulting in a weight vector denoted by $(W_{1},1-W_{1})$, where the superscript $(t)$ is excluded for simplicity. Moreover, assume $\phi_{1}=\phi_{2}=\phi$, which might occur when two algorithms deliver nearly identical results, with one slightly outperforming the other. If $\mathrm{cov}(W_{1},\phi)=\rho$, by definition it follows that $\mathrm{cov}(1-W_{1},\phi)=-\rho$; this implies that $\sum_{k=1}^{2}\mathbb{E}[\phi W_{k}]=\sum_{k=1}^{2}\mathbb{E}[\phi]\,\mathbb{E}[W_{k}]$. Note that this is merely a basic illustration, and the assumption is explored in greater depth in Appendix B via simulations and real data examples.
We highlight that if one prefers to avoid randomization, then one can set the value of $u^{(t)}$ to 00 and the coverage guarantee still holds. In addition, we note that if the weight vector takes the form $w^{(t)}_{k}=1$, and $w^{(t)}_{j}=0$ for all $j\neq k$, then our procedure reduces to the selection of the $k$-th model. A potential drawback of the procedure is that it may produce a union of intervals, even when the input sets are intervals. This does not occur if the intersection of the initial intervals is non-empty (Gasparin and Ramdas,, 2024).
Building on our earlier discussion, consider two canonical examples for the loss as follows: for regression tasks, the loss is quantified by the interval’s size, and for classification tasks, by the count of elements within the set. We define the loss function $\ell(\cdot)$ as
$\ell:=\ell(\mathcal{C})=\begin{cases}\text{Lebesgue measure}(\mathcal{C}),\quad&\mathrm{if}\leavevmode\nobreak\ \mathcal{Y}\subseteq\mathbb{R},\\ |\mathcal{C}|,\quad&\mathrm{if}\leavevmode\nobreak\ \mathcal{Y}=\{y_{1},\dots,y_{D}\},\end{cases}$ (6)
where $\mathcal{C}$ denotes any set, $|\mathcal{C}|$ represents the cardinality of $\mathcal{C}$, and $\{y_{1},\dots,y_{D}\}$ describes a discrete set with $D$ elements. Specifically, (6) varies when dealing with a regression or classification problem. In general, one could choose some nondecreasing function of the Lebesgue measure (or of the cardinality), but the identity seems to be a sensible canonical choice, so we stick to it in this paper. At each iteration, we observe losses of the experts $\ell^{(t)}=(\ell_{1}^{(t)},\dots,\ell_{K}^{(t)})\in\{\mathbb{R}^{+}\}^{K}$ and the cumulative loss for a given method after $t$ rounds is given by $L_{k}^{(t)}:=\ell^{(1)}_{k}+\dots+\ell^{(t)}_{k}$.
A notable property of the majority vote is that it generates sets that are never larger than twice the weighted average size of the initial sets. We present here a useful result, which will be used in the following.
Lemma 2.4.
Let $\mathcal{C}^{(t)}$ be the set defined in (3), $\ell_{M}^{(t)}=\ell(\mathcal{C}^{(t)})$ be the loss function defined in (6) associated to $\mathcal{C}^{(t)}$ and $w^{(t)}$ be any realization of $W^{(t)}$. Then,
$\ell_{M}^{(t)}=\ell(\mathcal{C}^{(t)})\leq 2\sum_{k=1}^{K}w_{k}^{(t)}\ell(\mathcal{C}_{k}^{(t)}).$ (7)
Proof.
We start with the discrete case $\mathcal{Y}=\{y_{1},\dots,y_{D}\}$. In particular, we have that
$\begin{split}\ell_{M}^{(t)}&=\sum_{d=1}^{D}\mathds{1}\left\{\sum_{k=1}^{K}w_{k}^{(t)}\mathds{1}\left\{y_{d}\in\mathcal{C}_{k}^{(t)}\right\}>\frac{1}{2}+\frac{u^{(t)}}{2}\right\}\leq\sum_{d=1}^{D}\mathds{1}\left\{\sum_{k=1}^{K}w_{k}^{(t)}\mathds{1}\left\{y_{d}\in\mathcal{C}_{k}^{(t)}\right\}>\frac{1}{2}\right\}\\ &\stackrel{{\scriptstyle(i)}}{{\leq}}\sum_{d=1}^{D}2\sum_{k=1}^{K}w_{k}^{(t)}\mathds{1}\left\{y_{d}\in\mathcal{C}_{k}^{(t)}\right\}=2\sum_{k=1}^{K}w_{k}^{(t)}\sum_{d=1}^{D}\mathds{1}\left\{y_{d}\in\mathcal{C}_{k}^{(t)}\right\}=2\sum_{k=1}^{K}w_{k}^{(t)}\ell(\mathcal{C}_{k}^{(t)}),\end{split}$
where $(i)$ holds due to the fact that $\mathds{1}\{x>1\}\leq x$, if $x\geq 0$. The proof in the case $\mathcal{Y}\subseteq\mathbb{R}$ follows the same strategy, it simply replaces the summation in $d$ by an integral and is given in Gasparin and Ramdas, (2024). ∎
From (7), it is now clear that our goal is to find a good weighting system in order to narrow down the size of $\mathcal{C}^{(t)}$, particularly we aim to assign a higher weight to experts that consistently yield smaller intervals compared to others.
In the following, we describe a general methodology to infer the weights assigned to the various experts over time. In particular, we start by describing the method under an i.i.d. setting and then extend the approach under a (possible) continuous distribution shift over time, as described in Gibbs and Candès, (2021).
3 Dynamic merging via exponential weighted majority
As illustrated in Section 1.1, our goal is to achieve a set that is as small as possible, particularly (7) asserts that this can be accomplished by learning a weight system that prioritizes methods yielding smaller sets. Furthermore, we want that the cumulative loss of our proposed method, indicated by $L_{M}^{(t)}:=\ell_{M}^{(1)}+\dots+\ell_{M}^{(t)}$, is not too large if compared to the cumulative loss incurred by the best expert up to time, indicated by
$L_{*}^{(t)}:=\min_{k\in[K]}L_{k}^{(t)}.$
A possible way to update our vector of weights is the Exponential Weight (ew) or Hedge Algorithm introduced in Freund and Schapire, (1997), which corresponds to Algortihm 1 with $\eta^{(t)}=\eta$, for all $t$. One way to evaluate the weighted majority method is the hedge loss $H^{(t)}=h^{(1)}+\dots+h^{(t)}$, where $h^{(t)}$ is the dot product $h^{(t)}=w^{(t)}\cdot\ell^{(t)}$. This can be interpreted as the expected loss achieved by randomly selecting the expert $k$ with probability $w_{k}^{(t)}$.
Data: $\mathcal{C}_{1}^{(t)},\dots,\mathcal{C}_{K}^{(t)}$ at each round $t$, initial learning rate $\eta^{(0)}\geq 0$
$w_{k}^{(1)}\leftarrow 1/K,L_{k}^{(0)}\leftarrow 0,$ $k=[K]$;
for rounds $t=1,\dots,T$ do
       $\mathcal{C}^{(t)}\leftarrow\left\{s\in\mathcal{S}:\sum_{k=1}^{K}w_{k}^{(t)}\mathds{1}\{s\in\mathcal{C}_{k}^{(t)}\}>\frac{1+u^{(t)}}{2}\right\}$;
       Receive loss $\ell_{k}^{(t)}$, update $L_{k}^{(t)}:=L_{k}^{(t-1)}+\ell_{k}^{(t)}$;
       Update learning rate $\eta^{(t)}$;
       $w_{k}^{(t+1)}\leftarrow\exp\{-\eta^{(t)}L_{k}^{(t)}\}/\sum_{j=1}^{K}\exp\{-\eta^{(t)}L_{j}^{(t)}\}$;
end for
Algorithm 1 Exponentially Weighted Majority Vote
The algorithm updates the weights according to the performance of the methods during the different rounds inversely proportional to the exponential of their size multiplied by a parameter $\eta$. This “learning rate” $\eta$ plays a crucial role; if $\eta$ approaches zero, the weights approach a uniform weighting, and if $\eta\to\infty$ the algorithm reduces to the Follow-the-Leader strategy, which puts all the weight on the method with the smallest loss so far, as explained in de Rooij et al., (2014). In certain scenarios, determining the appropriate value for this parameter may be difficult; indeed, we do not know if there is a clear method outperforming the others or any constant works well in some situations but not in others.
Our proposed solution is to employ the AdaHedge algorithm, which dynamically adjusts the learning parameter over time as follows. The weights at each round are assigned according to $w^{(t)}_{k}\propto\exp\{-\eta^{(t)}L^{(t-1)}_{k}\}$, $k\in[K]$, with
$\eta^{(t)}:=\frac{\ln K}{\delta^{(1)}+\dots+\delta^{(t-1)}},$
where $\delta^{(t)}=h^{(t)}-m^{(t)}$ is the difference between the hedge loss $h^{(t)}$ and the mix loss,
$m^{(t)}:=-\frac{1}{\eta^{(t)}}\ln\left(w^{(t)}\cdot\exp\{-\eta^{(t)}\ell^{(t)}\}\right).$
The AdaHedge algorithm provides an upper bound on regret, defined as the difference between the hedge loss and the loss of the best method up to time $t$, specifically:
$\mathcal{R}^{(t)}:=H^{(t)}-L_{*}^{(t)}.$
Consider $\ell_{+}^{(t)}=\max_{k}\ell_{k}^{(t)}$ and $\ell_{-}^{(t)}=\min_{k}\ell_{k}^{(t)}$ to represent the smallest and largest loss in round $t$, and let $L_{+}^{(t)}$ and $L_{-}^{(t)}$ represent their cumulative sum. Additionally, define $S^{(t)}=\max\{s^{(1)},\dots,s^{(t)}\}$ denote the largest loss range, where $s^{(t)}=l_{+}^{(t)}-l_{-}^{(t)}$, after $t$ trials. Suppose our objective is to find an upper bound for regret after an arbitrary number of rounds $T$. According to Theorem 8 in de Rooij et al., (2014), we have
$H^{(T)}\leq L_{*}^{(T)}+2\left(S^{(T)}\ln K\frac{(L_{+}^{(T)}-L_{*}^{(T)})(L_{*}^{(T)}-L_{-}^{(T)})}{L_{+}^{(T)}-L_{-}^{(T)}}\right)^{1/2}+S^{(T)}\left(\frac{16}{3}\ln K+2\right).$ (8)
It should be noted that this upper limit depends on the observed losses but not on the number of trials. This implies that the bound remains unchanged when additional rounds where all experts have the same loss are introduced. Another noteworthy feature of the method is its invariance to translations or scalings of the loss function. Specifically, the sequence of weights remains unchanged when the loss vectors $\ell^{(t)}$ are multiplied by a positive constant and (or) a quantity is added to them.
It is now possible to bound the loss of the sets produced by the majority vote method.
Theorem 3.1.
The loss of the weighted majority vote set over $T$ rounds can be bounded as
$L_{M}^{(T)}\leq 2L_{*}^{(T)}+4\left(S^{(T)}\ln K\frac{(L_{+}^{(T)}-L_{*}^{(T)})(L_{*}^{(T)}-L_{-}^{(T)})}{L_{+}^{(T)}-L_{-}^{(T)}}\right)^{1/2}+2S^{(T)}\left(\frac{16}{3}\ln K+2\right).$ (9)
Proof.
From Lemma 2.4 we have that $\ell_{M}^{(t)}\leq 2h^{(t)}$. This implies that $L_{M}^{(T)}=\ell_{M}^{(1)}+\dots+\ell_{M}^{(T)}\leq 2H^{(T)}$; combining this result with (8) gives the desired bound. ∎
The above expression is identical to the regret bound (8), except for an additional factor of 2.
In certain instances, the Lebesgue measure of the set cannot be directly employed because of situations where the width of some intervals can be unbounded. For example, in adaptive conformal prediction (Gibbs and Candès,, 2021), there is no guarantee that the value of $\alpha$ is greater than zero. A possible solution is to define the loss function as an increasing, nonnegative, bounded function of the measure of the set (so that the weight of a predictor does not get permanently set to zero if its $\alpha$ happens to be negative in some round). We will discuss this example in the following sections.
4 Conformal online model aggregation: i.i.d. setting
Code to reproduce all experiments can be found at https://github.com/matteogaspa/coma.
As explained at the beginning, a relatively open question in the conformal prediction literature is how to do model selection or aggregation using conformal prediction. Since all methods do provide the same coverage level of $1-\alpha$ under i.i.d. data, we use a working definition of the “best” predictor as the one whose expected set dimension is smaller.
Below, we provide one answer to this question in an online data setting, by employing the algorithms introduced in the previous section. When our dynamic merging algorithm is used (with constant, adaptive or another stepsize rule) in this conformal context, we title the method as “conformal online model aggregation”, or COMA for short. In the i.i.d. setting, it is often the case that our adaptive stepsize update rule results in the weight vector rapidly putting almost full (unit) weight on the best predictor, resulting effectively in online model “selection”, but experiments in non-i.i.d. settings suggest that the weight vector can meaningfully fluctuate between different predictors if they perform better in different periods.
We start by examining our proposed methods, in a scenario where data are i.i.d., this implies that each algorithm at each iteration provides an interval with the guarantee described in (1). To examine the practical behavior of COMA, we apply it to a real-world dataset. Specifically, the dataset comprises 75 0007500075\,00075 000 observations pertaining to Airbnb apartments in New York City. The response variable is represented by the logarithm of the nightly price, while the covariates include information about the apartment and its geographical location. The data has been partitioned into 75 rounds, each consisting of 1 00010001\,0001 000 observations and a single test point. At each iteration, $K$ split-conformal prediction intervals based on the data observed at round $t$ are constructed using various algorithms. In the first simulation, $K=4$ regression algorithms were employed: linear model, lasso, ridge, and random forest. In the second scenario $K=5$ and also a neural network was utilized.
Figure 1: Figure shows the performance of our conformal online model aggregation algorithm. We plot the hedge loss ($h_{t}$) obtained during various iterations with either a constant or adaptive learning rate scheme, along with the loss of the Random Forest algorithm (left), and the loss from both Random Forest and Neural Network (right). The series have been smoothed using a moving average $(5,\frac{1}{5})$. In both cases, COMA with the adaptive learning rate schedule (and with rate equal to 1) quickly learns that RF is the best method.
The weights were adjusted using the COMA method with either fixed or adaptive $\eta$ settings as previously mentioned. For Algorithm 1 with fixed learning parameter throughout the iterations, two different values were selected, $\eta=0.1$ and $\eta=1$. In Figure 1, the hedge loss during various iterations in both scenarios is presented, along with the loss obtained by the best performing method(s). It can be seen that in both cases, that the hedge loss of the COMA algorithm with $\eta=1$ and of the adaptive COMA stabilizes at the loss level obtained by the random forest after a certain number of iterations, which is the best method. On the other hand, if $\eta$ is set to 0.1 the loss exhibits a slower decrease, attributed to the lower value of the learning parameter. In Figure 6 and Figure 5 of Appendix A, we report the weights assumed by the different regression algorithms during certain iterations. It can be seen that the COMA with adaptive parameter, after a few iterations, concentrates all the mass in the random forest, while the COMA algorithm with fixed learning parameter is more conservative in assigning the weights. This fact can be explained by the plot in Figure 7 of Appendix A where it is possible to note that in both cases ($K=4$ and $K=5$) the learning parameter $\eta^{(t)}$ is always higher than the levels $\eta=1$ and $\eta=0.1$.
5 COMA under distribution shift
As described in Section 1, conformal prediction offers a general approach to convert the output of any prediction algorithm into a prediction set. Although its validity hinges on the assumption of exchangeability, this assumption is frequently violated in numerous real-world applications. In the energy market, for instance, consumer behavior can undergo significant shifts in response to new legislation, major global events, and financial market fluctuations. In general, where the distribution of the data shift over time we cannot expect to obtain reliable results using standard conformal inference. In recent years, several works have been proposed to address this issue; some examples are Gibbs and Candès, (2021) and Angelopoulos et al., 2024a . For the introduction of the methods, in order to avoid confusion, the discussion refers to situations involving the use of just one algorithm. The case involving multiple experts is addressed in Section 5.1 and Section 5.2.
In Gibbs and Candès, (2021), adaptive conformal inference (ACI) is introduced as a novel method for constructing prediction sets in an online manner, providing robustness to changes in the marginal distribution of the data. Their work is based on the update of the error level $\alpha$ to achieve the desired level of confidence. Indeed, given the non-stationary nature of the data-generating distribution, conventional results do not guarantee a $1-\alpha$ coverage. However, at each time $t$, an alternative value $\alpha^{(t)}$ might exist, allowing the attainment of the desired coverage.
The procedure described in Gibbs and Candès, (2021) recursively updates the error rate $\alpha$ as follows:
$\alpha^{(1)}=\alpha,$ (10)
$\alpha^{(t)}=\alpha^{(t-1)}+\gamma(\alpha-\phi^{(t-1)}),\quad t\geq 2,$ (11)
where $\phi^{(t)}=\mathds{1}\{Y_{t}\notin\mathcal{C}^{(t)}(\alpha^{(t)})\}$ is the sequence of miscoverage events setting the error to $\alpha^{(t)}$, while $\gamma>0$ represents a step size parameter. Possible improvements are proposed in Gibbs and Candès, (2023) and Zaffran et al., (2022), in particular, their proposals tune the parameter $\gamma$ over time using an online learning procedure. A drawback arises from the fact that in certain situations, ACI can return either $\mathbb{R}$ or $\emptyset$ as a set, attributed to the circumstance that in some iterations, the parameter $\alpha^{(t)}$ can be smaller than 0 or greater than 1.
Another method is presented in Angelopoulos et al., 2024a , where instead of updating the error level $\alpha$, there is an update of the quantile of the distribution of the scores. In this setting, we define $s^{(t)}:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ as the conformal score function, which measures the accuracy of the prediction at time $t$. Actually, if different models are involved one can use different scores (e.g. absolute or signed residuals), but in the following we will assume the same $s(\cdot,\cdot)$ for all algorithms. The procedure updates the quantiles as follows:
$q^{(1)}=q,$
$q^{(t)}=q^{(t-1)}+\gamma(\phi^{(t-1)}-\alpha),\quad t\geq 2,$
where $\phi^{(t)}=\mathds{1}\{y^{(t)}\notin\mathcal{C}^{(t)}(q^{(t)})\}$ and the sets at each time are simply defined by
$\mathcal{C}^{(t)}(q^{(t)})=\left\{y\in\mathcal{Y}:\,s^{(t)}(x^{(t)},y)\leq q^{(t)}\right\}.$
As noted in Angelopoulos et al., 2024a , both ACI and quantile tracking, can be recasted to an online descent algorithm with respect to the pinball loss (Koenker,, 2005). In addition, both methods achieve a long-run coverage in time, defined as
$\frac{1}{T}\sum_{t=1}^{T}\phi^{(t)}=\alpha+o(1),$ (12)
under few or no assumptions on $\{z^{(t)}\}_{t=1}^{T}$ and $\{s^{(t)}\}_{t=1}^{T}$, where $o(1)$ is quantity that tends to zero as $T\to\infty$. In particular, (15) holds deterministically and is different from the probabilistic assumption in (1) valid in an i.i.d. setting.
Also in the case of the quantile tracking some improvements are proposed, such as an adaptive learning rate or the introduction of an error integrator. The proposal in Angelopoulos et al., 2024b is to set a decaying step size $\gamma^{(t)}$. In particular, they prove that if scores are bounded and $\gamma^{(t)}\propto t^{-1/2-\epsilon}$ for some $\epsilon\in(0,1/2)$, then the miscoverage coverage error is bounded as $O(\frac{1}{T^{1/2-\epsilon}})$ in the adversarial setting, while in the i.i.d. the algorithm achieves a coverage guarantee.
In light of the dynamics shift observed in the marginal distribution of the data, it is conceivable that the weights assigned to different algorithms may also undergo variations throughout iterations. We now describe two different ways to combine our dynamic ensembling algorithm with ACI: either as a wrapper that does not alter the inner workings of ACI (Subsection 5.1), or by altering the feedback provided to ACI itself (Subsection 5.2).
5.1 Decentralized COMA under distribution shift
In the scenario described above, suppose that each agent operates independently by providing an interval at each iteration and, subsequently, the aggregator merges these intervals according to the vector $w^{(t)}$. This framework can be referred to as a decentralized dynamic merging. Under this setting, the advantage of the quantile tracking method is that it never returns sets of infinite dimensions. This implies that the loss function used can remain the size of the different sets, ensuring that the bound described in (9) is valid. In the case of using the described ACI approach, it becomes necessary to employ a function $g(\cdot)$ that is capable of returning a finite number when the miscoverage rate $\alpha_{k}^{(t)}$ is smaller than 0.
Before proceeding with the rest of the section, we introduce the following proposition, which holds under the following assumptions:
  1. 1.
    Negative total elementwise empirical correlation between $\{\phi^{(t)}\}_{t=1}^{T}$ and $\{w^{(t)}\}_{t=1}^{T}$:
    $(T^{-1})\sum_{t=1}^{T}\sum_{k=1}^{K}(\phi_{k}^{(t)}-\bar{\phi}_{k})(w_{k}^{(t)}-\bar{w}_{k})\leq 0,$ (13)
    where $\bar{\phi}_{k}:=(T^{-1})\sum_{t=1}^{T}\phi_{k}^{(t)}$ and $\bar{w}_{k}:=(T^{-1})\sum_{t=1}^{T}w_{k}^{(t)}$.
  2. 2.
    No randomization: the value $u^{(t)}$ in (3) equals 0, in symbols,
    $u^{(t)}=0,\quad t=1,\dots,T.$ (14)
Proposition 5.1.
Let $\{z^{(t)}=(x^{(t)},y^{(t)})\}_{t=1}^{T}$ be a sequence of data points, $\{w^{(t)}=(w_{1}^{(t)},\dots,w_{k}^{(t)})\}_{t=1}^{T}$ be a sequence of weight vectors and $\{\phi_{k}^{(t)}\}_{t=1}^{T}$ be the sequence of miscoverage events for the $k$-th algorithm, $k=1,\dots,K$. Under the assumptions in (13) and (14), then the decentralized dynamic merging in conjunction with adaptive conformal inference defined in (10) and (11) satisfies
$\frac{1}{T}\sum_{t=1}^{T}\mathds{1}\{y^{(t)}\notin\mathcal{C}^{(t)}\}\leq 2\alpha+2\frac{1-\alpha+\gamma}{T\gamma},$ (15)
for all $\alpha\in[0,1/2]$.
The proof is given in Appendix C, and similar results can be obtained using the quantile tracking method under the assumption of bounded scores (see Proposition 1 in Angelopoulos et al., 2024a and Theorem 1 in Angelopoulos et al., 2024b ). Randomization is not included because the UMI inequality is inapplicable in deterministic contexts however, the results using randomization appear to be reasonably decent also in this setting. The result (15) expresses a long-run coverage guarantee of level $1-2\alpha$ under arbitrary distribution if certain conditions regarding the relationship between the weights and the miscoverage events are fulfilled. Such conditions are challenging to verify in reality; however, empirically, the method appears to yield satisfactory results in real-world applications (see Appendix B for other experiments).
We now use the methods described in Section 3 in a real data set. We define $K=3$ different regression algorithms, and we employ quantile tracking with a decaying step size $\gamma^{(t)}\propto t^{-1/2-\epsilon}$. In particular, we use the ELEC2 (Harries,, 2003) data set that monitors electricity consumption and pricing in the states of New South Wales and Victoria in Australia, with data recorded every 30 minutes over a 2.5-year period from 1996 to 1999. For our experiment, we utilize four covariates: nswprice and vicprice, representing the electricity prices in each respective state, and nswdemand and vicdemand, denoting the usage demand in each state. The response variable is transfer, indicating the quantity of electricity transferred between the two states. We narrow our focus to a subset of the data, retaining only observations within the time range of 9:00 am to 12:00 pm to mitigate daily fluctuation effects, the same procedure is chosen in Barber et al., (2023). Additionally, we discard an initial time segment during which the transfer value remains constant. The models used correspond to three distinct linear models based on different covariates. The first model relies on the entire set of covariates at time $t$, the second model employs $y^{(t-1)}$ and $y^{(t-2)}$ as covariates (it corresponds to an AR(2) without constraints on the parameters), while the third model supplements the latter with the covariate nwsprice.
Figure 2: The top row displays local coverage for adaptive conformal inference using quantile tracking for the three regression algorithms described in Section 5.1. The bottom row shows local coverage for our merged set using COMA with adaptive $\eta$ or $\eta=0.1$ (with and without randomization). In the first row the dashed lines represent the level $1-\alpha$, while in the second column dashed lines represent the level $1-2\alpha$. In the bottom-left plot the lines of the randomized and non-randomized methods overlap as the weights concentrate after a few iterations onto a single model.
The performance of the methods is compared using the local level coverage on $200$ data points:
$\textrm{localCov}^{(t)}:=\sum_{i=t-100+1}^{t+100}\mathds{1}\left\{y^{(i)}\in\mathcal{C}_{k}^{(i)}\left(q_{k}^{(i)}\right)\right\},$ (16)
where $\mathcal{C}_{k}^{(t)}(q_{k}^{(t)})$ is the set obtained by the $k$-th regression algorithms with the quantile fixed at $q_{k}^{(t)}$. The local coverage is also used to measure the performance of our merging procedure. In particular, we compare the use of an adaptive learning parameter and a small fixed $\eta=0.1$. As can be seen in Figure 2, the COMA procedure with adaptive $\eta^{(t)}$ obtains a coverage that is similar to the prespecified level $1-\alpha$. This is because the weights concentrate rapidly on one model over time; in particular, this model is not always the same, as is possible from Figure 3. However, in both cases, the coverage is similar to the pre-specified level $1-\alpha$.
Figure 3: Weights (first row) and losses, $\ell^{(t)}$, (second row) obtained by the various model for the decentralized COMA algorithm applied in ELEC2 dataset (Section 5.1). The series of the losses have been smoothed using a moving average $(100,\frac{1}{100})$. The loss of the weighted majority vote stabilizes around the two best models after few iterations. The strategy with an adaptive $\eta^{(t)}$ is more aggressive than the one with $\eta^{(t)}=\eta$.
5.2 Adaptive conformal inference directly applied on dynamic merging
The approach described in the preceding section had agents updating the quantiles of their score function based on their own past errors. This makes sense when all agents can observe the ground truth and calculate their own errors, and they do not care about the aggregator’s goals. In other words, the dynamic merging algorithm was just an outer wrapper that did not interfere with the inner functioning of the $K$ adaptive conformal inference algorithms.
Below, we show that our dynamic ensembling method can provide direct feedback to the adaptive conformal inference framework, and this can have some advantages in settings where the end goal is good aggregator performance only, and individual agents do no have their own goals (of maintaining coverage, say). First, we observe that if the absolute residuals are employed as the score function and if the quantile ($q^{(t)})$ is common across all algorithms, then the size of the sets at time $t$ is equal for all intervals. This implies that the weights are not updated during the various iterations. In this case, we propose to use a value $\alpha^{(t)}$ that is common for all $K$ agents, and it is updated by the aggregator based on the performance of the exponential weighted majority vote. In other words, the aggregator, at each iteration, tries to learn the coverage level required to obtain a $(1-\alpha)$-confidence set. The procedure recursively updates the $\alpha$ level according to the miscoverage event $\phi^{(t)}=\mathds{1}\{Y_{t}\notin\mathcal{C}^{(t)}(\alpha^{(t)})\}$, where
$\mathcal{C}^{(t)}:=\left\{y\in\mathcal{Y}:\sum_{k=1}^{K}w_{k}^{(t)}\mathds{1}\left\{y\in\mathcal{C}_{k}^{(t)}\left(\alpha^{(t)}\right)\right\}>1/2\right\},$
and $w_{k}^{(t)}$ are the weights learned by the COMA procedure. The error level (common among the agents) is updated as in (10) and (11). It is important to remark that the length of the intervals can be different across the various agents also if $\alpha^{(t)}$ is common. In such a scenario, ACI operates directly on the set produced by the aggregator and so we directly recover the properties of the ACI method. Further, there is no need for randomization to enhance the length and coverage of the intervals. In particular, Gibbs and Candès, (2021) proves that the empirical error converges almost surely to $\alpha$ under weak conditions.
Figure 4: Local level coverage (first column), weights assumed during the iterations (second column) and losses (third column) for direct ACI for COMA with adaptive $\eta$ (top) and $\eta=0.1$ (bottom) applied in ELEC2 dataset (Section 5.2). The dashed line in the first column represents the level $1-\alpha$. The series of the losses have been smoothed using a moving average $(100,\frac{1}{100})$. In both cases, the local level coverage oscillates around the level $1-\alpha$.
We applied the procedure to the same dataset used previously, with the target coverage level fixed at 0.05 and the loss function is $\arctan(\ell^{(t)})$. The use of a function which bounds our loss is necessary since ACI can output $\mathbb{R}$ as interval. Also in this we use (16) to compare the various methods. The local level coverage oscillates around the prespecified while the it is possible to see that the strategy with an adaptive $\eta$ is more aggressive and adapts rapidly the weights according to the observed loss.
6 Summary
Our work presents a method for aggregating conformal prediction intervals derived from different algorithms. Specifically, a method based on majority voting is proposed, where the weight of each individual agent is updated based on its past performance. In some cases, when there is an algorithm that outperforms the others, the weight is assigned to the better-performing model, resulting in model selection. Particularly, in the case of i.i.d. data, it is very common for there to exist a best model that outperforms the others in terms of size of the outputted sets. It is feasible to establish an upper bound for the loss incurred by our method relative to the loss of the best-performing algorithm, translating into a form of regret for our method.
We also tested our method in cases where the data underwent a distribution shift, along with quantile tracking (Angelopoulos et al., 2024a, ; Angelopoulos et al., 2024b, ) and adaptive conformal inference (Gibbs and Candès,, 2021) methods. Specifically, two contexts are explored: in the first case, each model operates independently, while in the second case, the aggregator acts above the models, and adaptive conformal inference operates directly on it. Furthermore, even when the assumptions are not fully satisfied, the empirical results appear to meet the coverage expectations. Further research is required to assess whether our approaches can protect against a broader range of real-world situations.
Acknowledgments
AR acknowledges funding from NSF grant IIS-2229881, and thanks Rina Foygel Barber, Ryan Tibshirani and Emmanuel Candès for useful conversations around model selection in conformal prediction.
Appendix A Additional results from Section 4
Figure 5: Weights assumed by the regression algorithms during different iterations. The regression algorithms used are lasso, linear model, ridge and random forest.
Figure 6: Weights assumed by the regression algorithms during different iterations. The regression algorithms used are lasso, linear model, neural net, ridge and random forest.
Figure 7: Series of values taken by the learning parameter $\eta^{(t)}$ over time. The plot on the right corresponds to the plot on the left, zoomed in on the interval 0 to 10. The red dashed lines and the green dashed line represent the level 1 and 0.01. Solid lines denote the trajectory of the parameter $\eta^{(t)}$.
Appendix B Further experiments
B.1 Additional simulations on Adaptive Conformal Inference
Investigating the ACI (Adaptive Conformal Inference) method in conjunction with the adaptive COMA, we conduct a case study in which changes in marginal distribution also impact the weights of the two algorithms. We suppose to have two covariates, $x_{1}$ and $x_{2}$, and the data are generated in blocks, with only one of the two covariates influencing the response at a time. Specifically, at each time $t$ we have $y^{(t)}=2x^{(t)}+\epsilon^{(t)}$ where
$x^{(t)}=\begin{cases}x_{1}^{(t)},\quad t\in\{[0,50)\cup[100,150)\cup[250,350)\cup[450,550)\cup\dots\cup[1250,1350)\}\\ x_{2}^{(t)},\quad t\in\{[50,100)\cup[150,250)\cup[350,450)\cup\dots\cup[1350,1450]\}\end{cases}$
and $x_{1}^{(t)},x_{2}^{(t)},\epsilon^{(t)}$ are generated from a standard normal distribution. Two standard linear models are used, one using $x_{1}$ as a regressor and the other using $x_{2}$. Prediction intervals are generated using split conformal on the most recent 100 observations, and the parameter $\gamma$ for the Adaptive Conformal algorithm is set to 0.005. In addition, the cumulative density function of a $\textrm{Gamma}(1,0.1)$ is used to scale the length of the intervals. The entire procedure is repeated 2000 times. The choice of the weight assigned to each method dynamically adapts to the performance of the two approaches, evaluated in terms of cumulative loss, as can be seen in Figure 8. In particular, the weight assigned to one of the two experts is higher when the cumulative loss is lower than that of the other expert.
Figure 8: The mean of cumulative losses ($L_{1}^{(t)},L_{2}^{(t)}$) across 2000 replications throughout the iterations is depicted in the left plot for the linear model employing $x_{1}$ (orange) or $x_{2}$ (green) as the regressor. In the right plot, the average weights for 2000 replications are illustrated during the iterations of the linear model using either $x_{1}$ or $x_{2}$ as the regressor. In the background of both plots, the shaded region indicates iterations when $L_{1}^{(t)}>L_{2}^{(t)}$.
B.2 Amazon data
Inspired by Angelopoulos et al., 2024a , we analyze a historical series comprising the single stock open prices of Amazon from January 1, 2006, to December 31, 2014. The compared models consist of six different autoregressive models incorporating temporal lags from 1 to 6, trained on the logarithmic scale of the opening price, while the forecasts are obtained on the original scale. The method used to obtain the sets is the quantile tracking method described in Section 5 with $\alpha=0.1$, a similar analysis is performed in Angelopoulos et al., 2024a with an autoregressive model with 3 lags. The parameter $\gamma$ is updated as $\gamma^{(t)}=0.1B^{(t)}$, where $B^{(t)}$ is the maximum observed score in a trailing window of 100 iterations and the models are re-trained at every time step. The weights are updated using an adaptive $\eta$ or fixing $\eta$ to the level 0.01, this small value of $\eta$ allows the mixing of the different intervals during iterations.
Figure 9: Local level coverage obtained in a window of 100 data points (first column), weights assumed during the different iterations (second column), empirical sum over time of the quantity $\sum_{k=1}^{K}\phi_{k}^{(t)}w_{k}^{(t)}-\sum_{k=1}^{K}\phi_{k}^{(t)}\sum_{k=1}^{K}w_{k}^{(t)}$ obtained using an increasing number of observations (third column) and series of the original stock with corresponding prediction intervals (fourth column). The black dashed lines in the first column represent the level $1-2\alpha$. The result corresponding to the strategy with an adaptive $\eta$ are presented in the first row while the result corresponding to the strategy with $\eta=0.01$ are presented in the second row.
As shown in Figure 9, the local leverage coverage exceeds the level $1-2\alpha$ and approximates $1-\alpha$ in both instances. The approach using an adaptive $\eta$ aggressively allocates weights and it frequently performs model selection. Nonetheless, in each scenario, the AR(1) model outperforms others in the initial phase, whereas the AR(4) model appears more effective in the latter phase. The empirical sum of the quantity $\sum_{k=1}^{K}\phi_{k}^{(t)}w_{k}^{(t)}-\sum_{k=1}^{K}\phi_{k}^{(t)}\sum_{k=1}^{K}w_{k}^{(t)}$ it is around zero during all iterations in both cases.
B.3 Lasso regression with different tuning parameters
We create a synthetic dataset regarding a high-dimensional linear model in this way. We first simulate a matrix $X$ of dimension $3100\times 100$ from a multivariate normal with variance $I_{100}$, where $I_{a}$ denote an identity matrix of dimension $a\times a$. The outcome vector is defined as follows:
$y^{(t)}=\beta^{\mathrm{T}}x^{(t)}+0.3y^{(t-1)}+\varepsilon^{(t)}-0.3\varepsilon^{(t-1)},$
where $\beta$ is a coefficient vector with non-zero entries only in the first ten positions, $x^{(t)}$ represents the $t$-th row vector of $X$, and $\varepsilon^{(t)}$ is sampled from a standard Gaussian distribution. The data generation process follows an ARMA(1,1) model (see Chapter 2.6 in Tsay, (2005) for an introduction), augmented by a linear component that depends on the covariates in $X$. Here, the employed models are lasso regressions, each with varying penalty parameters $\lambda$. In particular, the $\lambda$ values are logarithmically equispaced in the interval $[-4,3]$. As the data are not exchangeable, the quantile tracking approach is utilized along with COMA to obtain prediction intervals. Throughout all iterations, the parameter $\gamma$ is set to 2.
In Figure 10, the best $\lambda$ oscillates between $\lambda_{4}$ and $\lambda_{5}$ for all iterations. The local level coverage is between the level $1-2\alpha$ and $1-\alpha$ in both cases. In the last column, it is possible to see that the difference between $\sum_{k=1}^{K}\phi_{k}^{(t)}w_{k}^{(t)}$ and $\sum_{k=1}^{K}\phi_{k}^{(t)}\sum_{k=1}^{K}w_{k}^{(t)}$ is around the zero level and decreases over time.
Figure 10: Local level coverage obtained in a window of 100 data points (first column), weights assumed during the different iterations (second column) and empirical sum over time of the quantity $\sum_{k=1}^{K}\phi_{k}^{(t)}w_{k}^{(t)}-\sum_{k=1}^{K}\phi_{k}^{(t)}\sum_{k=1}^{K}w_{k}^{(t)}$ obtained using an increasing number of observations (third column). The black dashed lines in the first column represent the level $1-2\alpha$. The result corresponding to the strategy with an adaptive $\eta$ are presented in the first row while the result corresponding to the strategy with $\eta=0.01$ are presented in the second row.
B.4 Additional simulations on the negative correlation assumption
In this section, we investigate, the assumption of negative correlation via simulation. Data are generated as follows:
$y^{(t)}=10\sin(\pi x_{1}^{(t)})+10\exp\{x_{2}^{(t)}\}-4x_{3}^{(t)}+0x_{4}^{(t)}+\varepsilon^{(t)},\quad t=1,\dots,300,$ (17)
where $x_{j}^{(t)}$, $j=1,\dots,4$, are generated from independent uniform random variable in (0,1) while $\varepsilon^{(t)}\sim\mathcal{N}(0,1)$. The burn-in period for training is equal to 100 observations and four different models are used: random forest, linear model, lasso, and neural net. The number of replications is equal to 2000 and all the results in Figure 11 are averaged over the number of replications.
Figure 11: Empirical correlation between $\phi_{k}^{(t)}$ and $w_{k}^{(t)}$ at every time step (first columns), empirical average of the weights (second column) and empirical calculation of the value $\sum_{k=1}^{K}(\mathbb{E}[W_{k}^{(t)}\phi_{k}^{(t)}]-\mathbb{E}[W_{k}^{(t)}]\mathbb{E}[\phi_{k}^{(t)}])$ (third column).
From Figure (11), it seems that the assumption of negative correlation is not always met for all iterations for each model, however, the empirical difference between $\sum_{k=1}^{K}\mathbb{E}[W_{k}^{(t)}\phi_{k}^{(t)}]$ and $\sum_{k=1}^{K}\mathbb{E}[W_{k}^{(t)}]\mathbb{E}[\phi_{k}^{(t)}]$ does not appear to be high. In this case, the neural net appears to be the best model, followed by the random forest. Indeed, this can be inferred from the non-linear relationship between the covariates and the response in (17).
B.5 Classification task in an online setting
In this section, we tackle the problem of classification within an online framework. A simulated dataset with eight distinct classes were generated, each class is originated from a bivariate normal distribution with mean vectors of $(i,i)$ and randomly generated covariance matrices for $i=1,\dots,8$. Additionally, two covariates were introduced to induce noise. We generated 250 data points per class, and treated the classification process as if the data points were arriving in a sequential manner. Specifically, following a burn-in phase of 1000 iterations, we retrained four different classifiers (LDA, QDA, Neural Net, Random Forest) at each step, and the weights for these algorithms were adjusted using the COMA method.
Figure 12: Local level coverage obtained in a window of 100 data points (first column), weights assumed during the different iterations (second column), empirical sum over time of the quantity $\sum_{k=1}^{K}\phi_{k}^{(t)}w_{k}^{(t)}-\sum_{k=1}^{K}\phi_{k}^{(t)}\sum_{k=1}^{K}w_{k}^{(t)}$ obtained using an increasing number of observations (third column) and series of the original stock with corresponding prediction intervals (fourth column). The black dashed lines in the first column represent the level $1-2\alpha$. The result corresponding to the strategy with an adaptive $\eta$ are presented in the first row while the result corresponding to the strategy with $\eta=0.01$ are presented in the second row.
The results are presented in Figure 12, where it is possible to note that the miscoverage lies between the level $1-\alpha$ and $1-2\alpha$. In this situation, the strategy with an adaptive $\eta$ gives very different weights if compared to the strategy with a $\eta=0.01$. In the last column, the sum of covariances is around zero for all iterations.
Appendix C Omitted proofs
Proof of Theorem 2.1.
By direct calculation, we can write
$\begin{split}\mathbb{P}\left(Y^{(t)}\notin\mathcal{C}^{(t)}\right)&=\mathbb{P}\left(\sum_{k=1}^{K}W_{k}^{(t)}\phi_{k}^{(t)}\geq\frac{1}{2}-U^{(t)}/2\right)\stackrel{{\scriptstyle(i)}}{{\leq}}2\mathbb{E}\left[\sum_{k=1}^{K}W_{k}^{(t)}\phi_{k}^{(t)}\right]=2\sum_{k=1}^{K}\mathbb{E}\left[W_{k}^{(t)}\phi_{k}^{(t)}\right]\\ \end{split}$
where $(i)$ is valid due to the uniformly-randomized Markov’s inequality described in Ramdas and Manole, (2023).
The first part of the theorem is a direct consequence of Hölder inequality. Indeed, due to the fact that $||W||_{1}=1$, we notice that
$\mathbb{E}\left[\sum_{k=1}^{K}\phi_{k}^{(t)}W_{k}^{(t)}\right]\leq\mathbb{E}\left[||\phi^{(t)}||_{\infty}||W^{(t)}||_{1}\right]=\mathbb{E}\left[||\phi^{(t)}||_{\infty}\right],$
where $\phi^{(t)}=(\phi_{1}^{(t)},\dots,\phi_{K}^{(t)})$.
The second part can be proved as follows: in the case of conditional coverage (case 1.), then it holds that
$\begin{split}\mathbb{P}\left(Y^{(t)}\notin\mathcal{C}^{(t)}\right)&\leq 2\sum_{k=1}^{K}\mathbb{E}\left[W_{k}^{(t)}\phi_{k}^{(t)}\right]=2\sum_{k=1}^{K}\mathbb{E}\left[\mathbb{E}\left[W_{k}^{(t)}\phi_{k}^{(t)}\mid\{Z^{(i)}\}_{i=1}^{T}\right]\right]\\ &=2\sum_{k=1}^{K}\mathbb{E}\left[W_{k}^{(t)}\mathbb{E}\left[\phi_{k}^{(t)}\mid\{Z^{(i)}\}_{i=1}^{T}\right]\right]\leq 2\alpha.\end{split}$
In the case of negative total elementwise correlation (case 2.) it holds that $\sum_{k=1}^{K}\mathbb{E}[W_{k}^{(t)}\phi_{k}^{(t)}]\leq\sum_{k=1}^{K}\mathbb{E}[W_{k}^{(t)}]\mathbb{E}[\phi_{k}^{(t)}]$ and under marginal coverage, we have that
$\mathbb{P}\left(Y^{(t)}\notin\mathcal{C}^{(t)}\right)\leq 2\sum_{k=1}^{K}\mathbb{E}\left[W_{k}^{(t)}\phi_{k}^{(t)}\right]\leq 2\sum_{k=1}^{K}\mathbb{E}[W_{k}^{(t)}]\,\mathbb{E}[\phi_{k}^{(t)}]\leq 2\alpha\sum_{k=1}^{K}\mathbb{E}[W_{k}^{(t)}]=2\alpha.$
This concludes the theorem. ∎
Proof of Proposition 5.1.
We remark that, if $u^{(t)}=0$, then $y^{(t)}$ is not contained in $\mathcal{C}^{(t)}$ if and only if $\sum_{k=1}^{K}w_{k}^{(t)}\mathds{1}\{y^{(t)}\notin\mathcal{C}_{k}^{(t)}\}\geq 1/2$. Then, since $\phi_{k}^{(t)}:=\mathds{1}\{y^{(t)}\notin\mathcal{C}_{k}^{(t)}\}$, we can write
$\begin{split}\sum_{t=1}^{T}\mathds{1}\{y^{(t)}\notin\mathcal{C}^{(t)}\}&=\sum_{t=1}^{T}\mathds{1}\left\{\sum_{k=1}^{K}w_{k}^{(t)}\mathds{1}\{y^{(t)}\notin\mathcal{C}_{k}^{(t)}\}\geq\frac{1}{2}\right\}\\ &\stackrel{{\scriptstyle(i)}}{{\leq}}\sum_{t=1}^{T}2\sum_{k=1}^{K}w_{k}^{(t)}\mathds{1}\{y^{(t)}\notin\mathcal{C}_{k}^{(t)}\},\\ &=2\sum_{k=1}^{K}\sum_{t=1}^{T}w_{k}^{(t)}\phi_{k}^{(t)}\\ &\stackrel{{\scriptstyle(ii)}}{{\leq}}2\sum_{k=1}^{K}\frac{1}{T}\left(\sum_{t=1}^{T}w_{k}^{(t)}\right)\left(\sum_{t=1}^{T}\phi_{k}^{(t)}\right)\\ &\stackrel{{\scriptstyle(iii)}}{{\leq}}2\sum_{k=1}^{K}\frac{1}{T}\left(\sum_{t=1}^{T}w_{k}^{(t)}\right)\left(T\alpha+\frac{1-\alpha+\gamma}{\gamma}\right)\\ &=2\left(\alpha+\frac{1-\alpha+\gamma}{T\gamma}\right)\sum_{k=1}^{K}\sum_{t=1}^{T}w_{k}^{(t)}\\ &=2T\alpha+2\frac{1-\alpha+\gamma}{\gamma},\\ \end{split}$
where $(i)$ holds due to the fact that $\mathds{1}\{x\geq 1\}\leq x,$ if $x\geq 0$ and $(ii)$ is due to assumption (13), in fact if the assumption holds, then $T\sum_{t=1}^{T}w_{k}^{(t)}\phi_{k}^{(t)}-\sum_{t=1}^{T}\phi_{k}^{(t)}\sum_{t=1}^{T}w_{k}^{(t)}\leq 0$. Proposition 4.1 in Gibbs and Candès, (2021) along with $\alpha\in[0,1/2]$ implies $(iii)$. Dividing both sides by $T$ gives the result. ∎